A company has just paid its first dividend of $4.01. Next year's dividend is forecast to grow by 10 percent, followed by another 10 per cent growth in year two. From year three onwards dividends are expected to grow by 3.3 percent per annum, indefinitely. Investors require a rate of return of 18 percent p.a. for investments of this type. The current price of the share is (round to nearest cent)
Answer #1

rate positively ..

computation of present value of Free cash flow
i ii iii=i+ii iv v vi=iv*v
year Dividend Terminal value total cash flow PVIF @ 18% present value
1     4.4110                4.41           0.8475              3.74
2     4.8521                 34.10              38.95           0.7182            27.97
Price today =            31.71
Terminal value = Dividend year 3/(required rate - growth rate)
4.8521*103.3%/(18%-3.3%)
           34.10
Answer =            31.71